# Global weak solutions for the inverse mean curvature flow in the Heisenberg group

### v2

@@ 1. Introduction
Moreover, if $B_{R_{0}}(g_{0})\subset\Omega^{c}\subset B_{\bar{R}}(g_{0})$ for some $0<R_{0}<2<\bar{R}$ then there exist $\widehat{C}>1$ depending only on $R_{0}$ and $\bar{R}$ and there exists $C_{0}>1$ depending only on $Q$, $\partial\Omega$ and $\bar{R}$, such that we have
(1.14) $(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{\bar{R}}-\log C_{0}\leq u^{\varepsilon}(g)\leq(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{R_{0}}+\varepsilon^{4}\log\widehat{C}\,,$
@@
for every $\varepsilon\in[0,1]$ and every $g\in\overline{\Omega}$. As a consequence, for any $\varepsilon\in[0,1]$ the solutions $u^{\varepsilon}$ are proper and their level sets $\Sigma_{s}:=\{u=s\}$ satisfty
@@
(1.15) $\Sigma_{s}\cap B_{\varphi_{\varepsilon}(s)}(g_{0})=\emptyset\,\qquad\hbox{and}\qquad\Sigma_{s}\subset B_{\psi(s)}(g_{0})\,\quad\hbox{for any}\quad s\geq 0\,,$
where ${\varphi_{\varepsilon}(s)=\frac{R_{0}}{\widehat{C}^{\varepsilon^{4}/Q-1}}\exp{\left(\frac{s}{Q-1}\right)}}$ and ${\psi(s)=C_{0}^{\frac{1}{Q-1}}\bar{R}\exp{\left(\frac{s}{Q-1}\right)}}$.
@@
The proof of the previous result follows from a standard compactness argument based on Corollary 1.3, passing to the limit the minimality property of the solutions $u^{\varepsilon}_{p}$ for the energy functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}(\cdot)$ in (1.7) as $p\to 1$ or $(\varepsilon,p)\to(0,1)$ respectively. Concerning the two-sided bound (1.14), the construction of subsolutions for Riemannian $p$-capacitary potential performed in Proposition 3.3 is precise enough to get the upper bound in (1.14) as $p\to 1$ in the whole range $\varepsilon\in[0,1]$. However, the lower bound in (1.14) and in turn properness of solutions turn out to be more subtle.
@@
In the sub-Riemannian case $\varepsilon=0$ properness can be readily inferred from the construction of barriers in terms of explicit $p$-capacitary potentials for each $p\in(1,Q)$ which give exact solutions of (1.6a)-(1.6b) in the complement of a Korányi ball $B_{\bar{R}}(g_{0})$ and in turn an exact solution as in (1.2) as $p\to 1$. On the other hand, when $\varepsilon>0$ no such explicit solution is known for $p>1$, therefore in the whole range $\varepsilon\in[0,1]$ we prefer to follow a different strategy, exploiting decay properties of the $p$-capacitary potentials $v^{\varepsilon}_{p}$ at infinity. Thus, we obtain the lower bound in (1.14) by combining for $q=\frac{Q(p-1)}{Q-p/2}$ an $L^{\infty}-L^{q}$ bound on annuli together with a global weak-$L^{\sigma}$ bound, for $\sigma=\frac{Q(p-1)}{Q-p}>q$, with controlled dependence on $p>1$ and uniformly in $\varepsilon\in(0,1]$. As discussed in more details in Section 5, this approach allows to handle the whole range $\varepsilon\in[0,1]$, the inequalities (1.14) being eventually a consequence of the pointwise decay $v^{\varepsilon}_{p}(g)\simeq\|g\|^{\frac{p-Q}{p-1}}$ as $\|g\|\to\infty$ for the $p$-capacitary potentials which is the natural counterpart of the weak-$L^{\sigma}$ integrability.
@@
The sharpness of the bounds in (1.14) and the related fact that the spheres in (1.15) do expand with the same speed, the latter depending on the homogeneous dimension $Q=2n+2$ of $\mathbb{H}^{n}$ in accordance with the sub-Riemannian nature of the space at infinity even for $\varepsilon\in(0,1]$, leave open the fundamental problem concerning the asymptotic behaviour of $u^{\varepsilon}$ as $\|g\|\to\infty$. In particular, it would be very interesting to know whether or not both the Riemannian and the sub-Riemannian IMCF evolution of the initial surface $\Sigma_{0}=\partial\Omega$ is asymptotically equivalent to a Korányi sphere with exponentially growing radius ${R_{s}\sim e^{\frac{s}{Q-1}}}$ as the time $s\to\infty$, in analogy with the results in [24] and [38] for the Euclidean case. The analysis of this problem seems to require a much better understanding of the fine asymptotic behaviour at infinity of the Riemannian $p$-capacitary potentials $v^{\varepsilon}_{p}$. In particular, both for a complement of a Korányi ball and for more general exterior domains, it would be interesting to study their decay properties as $\|g\|\to\infty$ in terms of an asymptotic expansion with an explicit dependence with respect to $p$ as $p\to 1$, in analogy with the analysis performed in [29] in the Euclidean spaces and in [4] on a class of manifolds with asymptotically nonnegative Ricci curvature.
@@
As the fine behaviour of $p$-capacitary potential on Riemannian manifold is gaining considerable interest in recent years, see, e.g., [22], [1], [2], [23], we find this aspect worth of further investigation. At the same time in a subsequent paper for the solution given in Theorem 1.4 we aim to investigate curvature bounds and higher regularity of the level sets $\left\{\Sigma^{\varepsilon}_{s}\right\}_{s\geq 0}\subseteq\mathbb{H}^{n}$, where $\Sigma^{\varepsilon}_{s}=\{u^{\varepsilon}=s\}$, together with the validity of the Geroch monotonicity formulas from [36] both in the Riemannian and the sub-Riemannian setting.
@@
The plan of the paper is as follows. In Section 2 we collect basic notations and facts about the Heisenberg group $\mathbb{H}^{n}$ and the relevant functional inequalities used here. In Section 3 we construct the relevant barriers for the boundary gradient estimate for $p$-capacitary potentials. In Section 4 we present two auxiliary results for $p$-harmonic functions, i.e., $L^{\infty}-L^{p}$ gradient estimates and Harnack inequality, which hold uniformly w.r.to the Riemannian approximation of the sub-Riemannian metric. In Section 5 we derive the weak-$L^{\sigma}$ bound and in turn the uniform pointwise two-sided bounds for the Riemannian $p$-capacitary potentials which hold globally in the domain. In Section 6 for the Riemannian $p$-capacitary potentials we derive gradient estimate at the boundary $\partial\Omega$ and we transfer it to the whole $\overline{\Omega}$ obtaining the Cheng-Yau’s inequality. In Section 7 we finally prove the main results of the paper, i.e., Theorem 1.2 and Theorem 1.4.
Acknowledgments: We thank Luca Capogna and Giovanna Citti for several interesting and useful discussions on the content of some results of Section 4 and the connections with their papers [9] and [10].
@@ 2. Basic notations and preliminary results
The Heisenberg group $\mathbb{H}^{n}$ is the non-abelian homogeneous Lie group $({\mathbb{R}}^{2n+1},\ast,\delta_{\lambda})$, where $\ast$ is the group operation and $\left(\delta_{\lambda}\right)_{\lambda>0}$ is the anisotropic family of dilations. More precisely, given two points $g=(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},t)$ and $g^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n},y^{\prime}_{1},\ldots,y^{\prime}_{n},t^{\prime})$ in $\mathbb{H}^{n}$,
@@
On the other hand at the identity $\mathfrak{h}_{1}=\mathbb{R}T$, where $T=\partial_{t}=[X_{i},Y_{i}]$, for $i=1,\ldots,n$, is the vertical vector field and coincides with the commutator of horizontal vector fields. More generally, we have $[X_{i},Y_{j}]=\delta_{i,j}T$ and $[X_{i},X_{j}]=[Y_{i},Y_{j}]=[X_{i},T]=[Y_{i},T]=0$ for $i,j=1,\ldots,n$, which will be useful below.
@@
$\nabla_{X_{i}}X_{j}$ $=\nabla^{\varepsilon}_{Y_{i}}Y_{j}=0,\quad\textrm{for all }i,j=1,\ldots,n,\quad\textrm{and}\quad\nabla^{\varepsilon}_{X_{i}}Y_{j}=0,\quad\textrm{for all }i\neq j,$
@@
whereas for $\varepsilon=0$ its horizontal divergence is given by ${\rm div}_{0}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]$
@@
[…]
@@
where ||$\cdot$|| ⋅ | is the Euclidean norm in $\mathbb{R}^{n}$ and to simplify the notation we denote from now on $(|x|^{2}+|y|^{2})$ simply by $|z|^{2}$, with $z=(x,y)$.
@@
so that $L_{g_{0}}(B_{r}(0))=B_{r}(g_{0})$ As it is well known, $d_{0}(\cdot,\cdot)$ and $\|(\,\cdot\,)^{-1}\cdot\|$ give equivalent distances and both induce the Euclidean topology on $\mathbb{H}^{n}$, so that only Korányi balls will be used for $\varepsilon=0$.
Here we recall that for an open set with smooth boundary $\Omega\subseteq\mathbb{H}^{n}$ a point $p\in\partial\Omega$ is characteristic if $T_{p}\Omega=span\{X_{1}(p),\ldots,X_{n}(p),Y_{1}(p),\ldots,Y_{n}(p)\}$. In case of a Korány ball radius $r>0$ centered at the origin the characteristic points are precisely $g=(0,0,\pm r^{2}/4)$.
@@
Clearly such an assumption amounts to require a quantitative one-sided flatness of $\partial\Omega$ at any of its characteristic point and it is clearly satisfied by the complement of any gauge-ball but also by gauge balls themselves (for the latter observe that for any $g\in\partial B_{r}(\bar{g})$ under the choice $g_{0}=g*{\bar{g}}^{-1}*g$ and $R_{0}=r$ it is straightforward to check that the ball $B_{R_{0}}(g_{0})$ has the desired properties because of Euclidean convexity of gauge balls).
Given the Riemannian metrics $|\cdot|_{\varepsilon}^{2}$, $\varepsilon>0$, the corresponding volume measures $dvol_{\varepsilon}(\cdot)$ are left-invariant and $dvol_{\varepsilon}(\cdot)=\varepsilon^{-1}\mathcal{L}^{2n+1}$, i.e., they coincide with the Lebesgue measure up to a constant factor. Note that, besides translation invariance, the Lebesgue measure under dilation satisfies the equality $\mathcal{L}^{2n+1}(\delta_{\lambda}(E))=\lambda^{Q}\mathcal{L}^{2n+1}(E)$ for any measurable subset $E\subseteq\mathbb{H}^{n}$ and for any $\lambda>0$. For bounded open set $\widetilde{\Omega}\subset\mathbb{H}^{n}$ with $C^{1}$-smooth boundary and $\varepsilon>0$ we will also consider the notion of Riemannian perimeter of $\widetilde{\Omega}$ according to the standard definition, by duality with vector fields $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{H}^{n}\to\mathbb{R}^{2n+1}$, given by
As it is well known, due to the Riemannian divergence theorem we have $Per_{\varepsilon}(\widetilde{\Omega})=\mathcal{H}^{2n}_{\varepsilon}(\partial\widetilde{\Omega})$, where $\mathcal{H}^{2n}_{\varepsilon}(\cdot)$ is the Hausdorff measure associated to $d_{\varepsilon}$, which is also equal to the area computed with respect to the Riemannian volume form coming from the metric induced on $\partial\widetilde{\Omega}$. In the same way for $\varepsilon=0$ following, e.g., [12], [33] and [21], we will consider the orizontal perimeter $Per_{0}(\widetilde{\Omega})$ computed with a similar formula involving the horizontal divergence, horizontal vector fields (i.e., those with $\varphi_{2n+1}\equiv 0$) and where integration is with respect to $\mathcal{L}^{2n+1}$. The useful convergence property $\varepsilon Per_{\varepsilon}(\widetilde{\Omega})\to Per_{0}(\widetilde{\Omega})$ as $\varepsilon\to 0$ will be be discussed in Lemma 5.2 together with a representation formula for any $\varepsilon\geq 0$ in the spirit of [12, Section 3] and [33, Section 5].
Next, we recall three basic properties of the spaces $(\mathbb{H}^{n},d_{\varepsilon},dvol_{\varepsilon})$ which will be relevant in the sequel. Measures $dvol_{\varepsilon}(\cdot)$ are uniformly doubling, i.e., according to [19, Theorem 3.6], there exists constant $C_{D}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$ we have
@@
It follows from the doubling property that $dvol_{\varepsilon}(\cdot)$ support weak $(1,1)$-Poincarè inequality with uniform constants. More precisely, according to [19, Theorem 4.2], there exist constants $C_{P}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$ we have
@@
It is a remarkable consequence of (2.2) and (2.3) that Sobolev inequality also holds in a localized version uniformly on $\varepsilon>0$ (see [19, Theorem 4.6], see also [9, Lemma 7.4]). Thus, for any $1<p<Q$ there exists a constant $C_{S}$ (depending only on $p$ and $Q$) such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$
@@
The proofs of (2.2), (2.3), and (2.5) in [19] rely on the equivalence (uniformly on $\varepsilon>0$, see [19, Corollary 3.9] ) of the distance $d_{\varepsilon}$ with the $\varepsilon$-gauge $\mathcal{N}_{\varepsilon}$, where $\mathcal{N}_{\varepsilon}(z,t)=|z|+{\rm min}\left\{\frac{|t|}{\varepsilon},\sqrt{|t|}\right\}$ (see also [9] for an alternative definition of equivalent $\varepsilon$-gauge). Such equivalence makes more transparent, at least on the subgroup $\mathbb{R}\simeq\{(0,0,t)\in\mathbb{H}^{n},\,\,t\in\mathbb{R}\}$, the Euclidean character of $d_{\varepsilon}$ at distances of order below $\varepsilon^{2}$ and its sub-Riemannian character otherwise. Here we adopt still another definition of a (uniformly) equivalent $\varepsilon$-gauge, namely $\|\cdot\|_{\varepsilon}$, where
@@
so that $\|\cdot\|_{\varepsilon}\to\|\cdot\|$ pointwise on $\mathbb{H}^{n}$ as $\varepsilon\to 0$. Then, straightforward computations in combination with [19, Corollary 3.7 and Corollary 3.9] lead to the following result that collects relevant properties the $\varepsilon$-gauge above and their consequences.

### v3

@@ 1. Introduction
Moreover, if $B_{R_{\ast}}(g_{0})\subset\Omega^{c}\subset B_{\bar{R}}(g_{0})$ for some $0<R_{\ast}<2<\bar{R}$ then there exist $\widehat{C}>1$ depending only on $R_{\ast}$ and $\bar{R}$ and there exists $C_{0}>1$ depending only on $Q$, $R_{\ast}$, $_{0}, R¯\bar{R}$ and $\partial\Omega$, such that we have
(1.14) $(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{\bar{R}}-\log C_{0}\leq u^{\varepsilon}(g)\leq(Q-1)\log\frac{\|g_{0}^{-1}*g\|}{R_{\ast}}+\varepsilon^{4}\log\widehat{C}\,$,
@@
for every $\varepsilon\in[0,1]$ and every $g\in\overline{\Omega}$. As a consequence, for any $\varepsilon\in[0,1]$ the solutions $u^{\varepsilon}$ are proper and their level sets $\Sigma^{\varepsilon}_{s}:=\{u^{\varepsilon}=s\}$ satisfy
@@
(1.15) $\Sigma^{\varepsilon}_{s}\cap B_{\varphi_{\varepsilon}(s)}(g_{0})=\emptyset\,\qquad\hbox{and}\qquad\Sigma^{\varepsilon}_{s}\subset B_{\psi(s)}(g_{0})\,\quad\hbox{for any}\quad s\geq 0\,$,
where ${\varphi_{\varepsilon}(s)=\frac{R_{\ast}}{\widehat{C}^{\varepsilon^{4}/Q-1}}\exp{\left(\frac{s}{Q-1}\right)}}$ and ${\psi(s)=C_{0}^{\frac{1}{Q-1}}\bar{R}\exp{\left(\frac{s}{Q-1}\right)}}$.
@@
The proof of the previous result follows from a standard compactness argument based on Corollary 1.3, passing to the limit the minimality property of the solutions $u^{\varepsilon}_{p}$ for the energy functionals $J^{p,\varepsilon}_{u^{\varepsilon}_{p}}(\cdot)$ in (1.7) as $p\to 1$ or $(\varepsilon,p)\to(0,1)$ respectively. Concerning the two-sided bound (1.14), the construction of subsolutions for Riemannian $p$-capacitary potential performed in Proposition 3.3 is precise enough to get the upper bound in (1.14) as $p\to 1$ in the whole range $\varepsilon\in[0,1]$ by an iterative argument of self-improving nature on expanding annuli. However, the lower bound in (1.14) and in turn properness of solutions turn out to be more subtle, as we discuss below.
@@
In the sub-Riemannian case $\varepsilon=0$ properness can be readily inferred from the construction of barriers in terms of explicit $p$-capacitary potentials for each $p\in(1,Q)$ which give exact solutions of (1.6a)-(1.6b) in the complement of a Korányi ball $B_{\bar{R}}(g_{0})$ and in turn an exact solution as in (1.2) as $p\to 1$. On the other hand, when $\varepsilon>0$ no such explicit solution is known for $p>1$, therefore in the whole range $\varepsilon\in[0,1]$ we have to follow a different strategy, exploiting decay properties of the $p$-capacitary potentials $v^{\varepsilon}_{p}$ at infinity. Thus, we obtain the lower bound in (1.14) by combining for $q=\frac{Q(p-1)}{Q-p/2}$ an $L^{\infty}-L^{q}$ bound on annuli together with a global weak-$L^{\sigma}$ bound, for $\sigma=\frac{Q(p-1)}{Q-p}>q$, with controlled dependence on $p>1$ and uniformly in $\varepsilon\in(0,1]$. As discussed in more details in Section 5, this approach allows to handle the whole range $\varepsilon\in[0,1]$, the inequalities (1.14) being eventually a consequence of the pointwise decay $v^{\varepsilon}_{p}(g)\simeq\|g\|^{\frac{p-Q}{p-1}}$ as $\|g\|\to\infty$ for the $p$-capacitary potentials which is the natural counterpart of the weak-$L^{\sigma}$ integrability.
@@
The bounds in (1.14) being sharp, they show that the weak IMCF is trapped between spheres that in view of (1.15) do expand with the same speed depending only on the homogeneous dimension $Q=2n+2$ of $\mathbb{H}^{n}$, which is striking consequence of the sub-Riemannian nature of the space at infinity even for $\varepsilon\in(0,1]$. Of course the aforementioned bounds leave open the fundamental problem concerning the asymptotic behaviour of $u^{\varepsilon}$ as $\|g\|\to\infty$. In particular, it would be very interesting to know whether or not both the Riemannian and the sub-Riemannian IMCF evolution of the initial surface $\Sigma_{0}=\partial\Omega$ is asymptotically equivalent to a Korányi sphere with exponentially growing radius ${R(s)\sim e^{\frac{s}{Q-1}}}$ as the time $s\to\infty$ (i.e., whether or not once area-normalized under group dilations the surfaces converge to a Korányi sphere with fixed radius), in analogy with the results in [25] and [41] for the Euclidean case.
@@
A deeper analysis of the large time behaviour of the flow seems to require a much better understanding of the fine asymptotic behaviour at infinity of the Riemannian $p$-capacitary potentials $v^{\varepsilon}_{p}$. In particular, both for a complement of a Korányi ball and for more general exterior domains, it would be interesting to study their decay properties as $\|g\|\to\infty$ in terms of an asymptotic expansion with an explicit dependence with respect to $p$ as $p\to 1$, in analogy with the analysis performed in [30] in the Euclidean spaces and in [4] on a class of manifolds with asymptotically nonnegative Ricci curvature, at least for fixed $p>1$. As the fine behaviour of $p$-capacitary potential on Riemannian manifold is gaining considerable interest in recent years, see, e.g., [23], [1], [2], [24] and [33], we find this aspect worth of further investigation in the case of the Heisenberg group $\mathbb{H}^{n}$. At the same time in a subsequent paper for the solution given in Theorem 1.4 we aim to investigate curvature bounds and higher regularity of the level sets $\left\{\Sigma^{\varepsilon}_{s}\right\}_{s\geq 0}\subseteq\mathbb{H}^{n}$, with $\Sigma^{\varepsilon}_{s}=\{u^{\varepsilon}=s\}$ to be considered as curvature varifolds, together with the validity of the Geroch monotonicity formulas from [38] both in the Riemannian and the sub-Riemannian setting.
@@
The plan of the paper is as follows. In Section 2 we collect basic notations and facts about the Heisenberg group $\mathbb{H}^{n}$ and the relevant functional inequalities used here. In Section 3 we construct the relevant barriers for the boundary gradient estimates for $p$-capacitary potentials. In Section 4 we present two auxiliary results for $p$-harmonic functions, i.e., $L^{\infty}-L^{p}$ gradient estimates and Harnack inequality, which hold uniformly w.r.to the Riemannian approximation of the sub-Riemannian metric. In Section 5 we derive the weak-$L^{\sigma}$ bound and in turn the uniform pointwise two-sided bounds for the Riemannian $p$-capacitary potentials which hold globally in the domain. In Section 6 we derive the key Bochner inequality associated to linearized $p$-Laplacian in turn based on a refined Kato-type inequality for the rough Hessian. In Section 7 for the Riemannian $p$-capacitary potentials we derive gradient estimate at the boundary $\partial\Omega$ and we transfer it to the whole $\overline{\Omega}$ obtaining the Cheng-Yau’s inequality under a smallness assumption of the gradient at infinity. In Section 8 we finally prove the main results of the paper, i.e., Theorem 1.2 and Theorem 1.4.
@@ 2. Basic notations and preliminary results
The Heisenberg group $\mathbb{H}^{n}$ is the non-abelian homogeneous Lie group $({\mathbb{R}}^{2n+1},\ast,\delta_{\lambda})$, where $\ast$ is the group operation and $\left(\delta_{\lambda}\right)_{\lambda>0}$ is the anisotropic family of dilations (group homomorphisms). More precisely, given two points $g=(x_{1},\ldots,x_{n},y_{1},\ldots,y_{n},t)$ and $g^{\prime}=(x^{\prime}_{1},\ldots,x^{\prime}_{n},y^{\prime}_{1},\ldots,y^{\prime}_{n},t^{\prime})$ in $\mathbb{H}^{n}$,
@@
On the other hand at the identity $\mathfrak{h}_{1}=\mathbb{R}T$, where $T=\partial_{t}=[X_{i},Y_{i}]$, for $i=1,\ldots,n$, is the vertical vector field and coincides with the commutator of horizontal vector fields. More generally, we have $X_{i},Y_{j}]=\delta_{i,j}T$ and $T,T]=[X_{i},X_{j}]=[Y_{i},Y_{j}]=[X_{i},T]=[Y_{i},T]=0$ for $i,j=1,\ldots,n$, which will be useful below.
@@
$\nabla^{\varepsilon}_{T_{\varepsilon}}T_{\varepsilon}=0\,$$\quad$ $\nabla^{\varepsilon}_{X_{i}}X_{j}=\nabla^{\varepsilon}_{Y_{i}}Y_{j}=0\quad\textrm{for all }i,j=1,\ldots,n,\quad\textrm{and}\quad\nabla^{\varepsilon}_{X_{i}}Y_{j}=\nabla^{\varepsilon}_{Y_{i}}X_{j}=0,\quad\textrm{for all }i\neq j,$
@@
Notice in particular that for any fixed $j=1,\ldots,2n$ and for $E_{j}\in\{X_{1},\ldots,X_{n},Y_{1},\ldots,Y_{n}\}$ we have
(2.1) $\sum_{i=1}^{2n+1}|\nabla^{\varepsilon}_{E_{i}}E_{j}|_{\varepsilon}^{2}=\frac{1}{2\varepsilon^{2}}\qquad\hbox{whereas}\qquad\sum_{i=1}^{2n+1}|\nabla^{\varepsilon}_{E_{i}}T_{\varepsilon}|_{\varepsilon}^{2}=\frac{n}{2\varepsilon^{2}}\,.$
On the other hand, for $i,j,k\in\{1,\ldots,2n+1\}$ we have also the identities
(2.2) $\langle\nabla^{\varepsilon}_{E_{i}}E_{j},E_{k}\rangle_{\varepsilon}=-\langle E_{j},\nabla^{\varepsilon}_{E_{i}}E_{k}\rangle_{\varepsilon}\qquad\hbox{and}\qquad\langle\nabla^{\varepsilon}_{E_{i}}E_{j},\nabla^{\varepsilon}_{E_{i}}E_{k}\rangle_{\varepsilon}=0\quad\hbox{for}\quad j\neq k\,.$
@@
whereas for $\varepsilon=0$ its horizontal divergence is given by ${\rm div}_{0}Z_{\varphi}=\sum_{i=1}^{n}\left[(X_{i}\varphi_{i}+Y_{i}\varphi_{n+i}\right]$ .
@@
In terms of $|\cdot|_{\varepsilon}$ for $\varepsilon>0$ the Riemannian distance $d_{\varepsilon}$ between points $g,g^{\prime}\in\mathbb{H}^{n}$ is defined by minimizing $\ell(\gamma)=\int_{0}^{1}|\gamma^{\prime}(s)|_{\varepsilon}\,ds$ among absolutely continuous paths $\gamma:[0,1]\to\mathbb{H}^{n}$ joining $g$ and $g^{\prime}$, with the further restriction when defining the sub-Riemannian distance $d_{0}$ that $\gamma^{\prime}$ is a.e. horizontal. Note that $d_{\varepsilon}\leq d_{0}$ and indeed $d_{\varepsilon}\uparrow d_{0}$ as $\varepsilon\to 0$. For $\varepsilon>0$ we denote by $B^{\varepsilon}_{r}(g_{0})$ the metric balls $B^{\varepsilon}_{r}(g_{0})=\{g\in\mathbb{H}^{n}\,|\,\,d_{\varepsilon}(g_{0},g)<r\}$, or sometimes simply $B_{r}^{\varepsilon}$ the balls with $g_{0}=0$. Note that $B^{\varepsilon}_{r}(g_{0})=L_{g_{0}}(B^{\varepsilon}_{r}(0))$ and $\delta_{\lambda}(B^{\varepsilon}_{r}(g_{0}))=B^{\varepsilon\lambda}_{r\lambda}(\delta_{\lambda}(g_{0}))$ for any $g_{0}\in\mathbb{H}^{n}$ and for any $r>0$ and $\lambda>0$.
@@
where $|\cdot|$ is the Euclidean norm in $\mathbb{R}^{n}$ and to simplify the notation we denote from now on $(|x|^{2}+|y|^{2})$ simply by $|z|^{2}$, with $z=(x,y)\in\mathbb{R}^{2n}$.
@@
so that $L_{\hat{g}}(B_{r}(g_{0}))=B_{r}(\hat{g}*g_{0})$ and $\delta_{\lambda}(B_{r}(g_{0}))=B_{r\lambda}(\delta_{\lambda}(g_{0}))$ for any $\hat{g}\in\mathbb{H}^{n}$ and any $\lambda>0$. As it is well known, $d_{0}(\cdot,\cdot)$ and $\|(\,\cdot\,)^{-1}\cdot\|$ give equivalent distances and both induce the Euclidean topology on $\mathbb{H}^{n}$, so that only Korányi balls will be used for $\varepsilon=0$.
Here we recall that for an open set with smooth boundary $\Omega\subseteq\mathbb{H}^{n}$ a point $p\in\partial\Omega$ is characteristic if $T_{p}\Omega=span\{X_{1}(p),\ldots,X_{n}(p),Y_{1}(p),\ldots,Y_{n}(p)\}$. In case of a Korány ball $B_{r}(0)$ it is straightforward to check that the characteristic points are precisely $g=(0,0,\pm r^{2}/4)$.
@@
We say that an open set $\Omega\subseteq\mathbb{H}^{n}$ satisfy the exterior uniform gauge-ball condition ($\mathbf{HP_{\Omega}}$) with parameter $R_{0}>0$ if there exists $R_{0}>0$ such that for any $g\in\partial\Omega$ there exists $g_{0}\in\mathbb{H}^{n}$ such that $B_{R_{0}}(g_{0})\subset\Omega^{c}$ and $g\in\partial\Omega\cap\partial B_{R_{0}}(g_{0})$, where $B_{r}(g_{0})$ is a gauge-ball associated with the Korányi norm $\|\cdot\|$.
Clearly for open sets with at least $C^{2}$-smooth boundary such an assumption amounts to require a quantitative one-sided flatness of $\partial\Omega$ at any of its characteristic point and it is clearly satisfied by the complement of any gauge-ball but also by gauge balls themselves (for the latter observe that for any $g\in\partial B_{r}(\bar{g})$ under the choice $g_{0}=g*{\bar{g}}^{-1}*g$ and $R_{0}=r$ it is straightforward to check that the ball $B_{R_{0}}(g_{0})$ has the desired properties because of Euclidean convexity of gauge balls).
Given the Riemannian metrics $|\cdot|_{\varepsilon}^{2}$, $\varepsilon>0$, the corresponding volume measures $dvol_{\varepsilon}(\cdot)$ are left-invariant and $dvol_{\varepsilon}(\cdot)=\varepsilon^{-1}\mathcal{L}^{2n+1}$, i.e., they coincide with the Lebesgue measure up to a constant factor. Note that, besides translation invariance with respect to the group law, the Lebesgue measure under dilation satisfies the equality $\mathcal{L}^{2n+1}(\delta_{\lambda}(E))=\lambda^{Q}\mathcal{L}^{2n+1}(E)$ for any measurable subset $E\subseteq\mathbb{H}^{n}$ and for any $\lambda>0$. For bounded open set $\widetilde{\Omega}\subset\mathbb{H}^{n}$ with $C^{1}$-smooth boundary and $\varepsilon>0$ we will also consider the notion of Riemannian perimeter of $\widetilde{\Omega}$ according to the standard definition, by duality with vector fields $Z_{\varphi}=\sum_{i=1}^{n}\left[\varphi_{i}X_{i}+\varphi_{n+i}Y_{i}\right]+\varphi_{2n+1}T_{\varepsilon}$ associated to a smooth map $\varphi:\mathbb{H}^{n}\to\mathbb{R}^{2n+1}$, given by
@@
Next, we recall three basic properties of the spaces $(\mathbb{H}^{n},d_{\varepsilon},dvol_{\varepsilon})$ which will be relevant in the sequel. Measures $dvol_{\varepsilon}(\cdot)$ are uniformly doubling, i.e., according to [20, Theorem 3.6], there exists constant $C_{D}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$ and $\bar{g}\in\mathbb{H}^{n}$ we have
@@
It follows from the doubling property that $dvol_{\varepsilon}(\cdot)$ support weak $(1,1)$-Poincarè inequality with uniform constants. More precisely, according to [20, Theorem 4.2], there exist constants $C_{P}\geq 1$ depending only on $Q$ such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$ we have
@@
It is a remarkable consequence of (2.4) and (2.5) that Sobolev inequality also holds in a localized version uniformly on $\varepsilon>0$ (see [20, Theorem 4.6], see also [9, Lemma 7.4]). Thus, for any $1<p<Q$ there exists a constant $C_{S}$ (depending only on $p$ and $Q$) such that for every $\varepsilon>0$, $\bar{g}\in\mathbb{H}^{n}$ and $r>0$
@@
The proofs of (2.4), (2.5), and (2.7) in [20] rely on the equivalence (uniformly on $\varepsilon>0$, see [20, Corollary 3.9] ) of the distance $d_{\varepsilon}$ with the $\varepsilon$-gauge $\mathcal{N}_{\varepsilon}$, where $\mathcal{N}_{\varepsilon}(z,t)=|z|+{\rm min}\left\{\frac{|t|}{\varepsilon},\sqrt{|t|}\right\}$ (see also [9] for an alternative definition of equivalent $\varepsilon$-gauge). Such equivalence makes more transparent, at least on the subgroup $\mathbb{R}\simeq\{(0,0,t)\in\mathbb{H}^{n},\,\,t\in\mathbb{R}\}$, the Euclidean character of $d_{\varepsilon}$ at distances of order below $\varepsilon$ (i.e., for $|t_{2}-t_{1}|=O(\varepsilon^{2})$) and its sub-Riemannian character otherwise. Here we adopt still another definition of a (uniformly) equivalent $\varepsilon$-gauge, namely $\|\cdot\|_{\varepsilon}$, where
@@
so that $\|\cdot\|_{\varepsilon}\to\|\cdot\|$ pointwise on $\mathbb{H}^{n}$ as $\varepsilon\to 0$. Then, straightforward computations in combination with [20, Corollary 3.7 and Corollary 3.9] lead to the following result that collects relevant properties the $\varepsilon$-gauge defined just above and their consequences.